OpenStudy (anonymous):

The center of a circle is located at (3, 8), and the circle has a radius that is 5 units long. What is the general form of the equation for the circle?

OpenStudy (anonymous):

A)x2 + y2 – 6x – 16y + 48 = 0 B)x2 + y2 – 6x – 16y – 25 = 0 C)x2 + y2 + 6x + 16y + 48 = 0 D)x2 + y2 + 6x + 16y – 25 = 0